there are 600 people residing in a campus. out of these 340 people take Hindustan Times news paper and 270 people take the Times of India. There are 70 people who take both the papers. Find the number of persons who do not take either of the two papers..//
1
Expert's answer
2011-12-29T10:13:17-0500
The number of persons who reads some paper is 270 + 340 - 70 = 540.
Here we subtrackted 70 cause sum 270 + 340 includes persons who takes both papers twice.
So, the number of persons who do not take either of the two papers is
N = 600 - 540 = 60.